# Codes Correcting Transpositions of Consecutive Symbols

### v1

@@ 3 Codes correcting $t=\textnormal{const}$ transpositions
The bound $M^{\star}_{2}(n;t)\gtrsim c_{t}\frac{2^{n+t}}{n^{t}}$ was derived in [16, Theorem 11]333[16, Theorem 11] states a slightly worse constant $c_{2}=\frac{1}{4}$ than Theorem 3.1 of the present paper. It can be shown based on [27, Theorem 3] that $c_{2}$ can be improved to $\frac{1}{3}$. by exhibiting a geometric interpretation of binary codes correcting successive transposition (a.k.a. bit-shift) errors as packings in the Manhattan metric, and using the densest known packings. Unfortunately, the mentioned interpretation fails when $q>2$ and thus the construction that implies this lower bound cannot be generalized to larger alphabets in any obvious way. $\blacklozenge$
4 Codes correcting $t=\tau n$ transpositions
In this section we consider the regime where the number of transpositions grows linearly with the block-length. We derive a lower bound on the asymptotic rate of optimal codes in $\mathbb{Z}_{q}^{n}$ correcting $t=\tau n$ transpositions, for any fixed $\tau\in[0,\frac{1}{2}]$ (Theorem 4.1 ahead). The bound will in fact be expressed as a function of the minimum distance, as is more common in the literature, but it can easily be written in terms of the number of correctable transpositions as well. Namely, Proposition 1 guarantees that codes of relative distance $\delta$ can correct a fraction of $\tau=\frac{\delta}{2}$ transpositions.
4.1 Generalized Gilbert–Varshamov bound
The standard Gilbert–Varshamov bound states that the cardinality of an optimal code of minimum distance $d$ is lower-bounded by the ratio of the cardinality of the code space and the cardinality of a ball of radius d1d-1 [8, 28]. The bound was later generalized [13, 10] to include nonuniform spaces in which the cardinality of a ball of given radius depends on its center. Namely, it was proven in [10] that the cardinality of an optimal code of minimum distance $d$ is lower-bounded by the ratio of the cardinality of the code space and the average cardinality of a ball of radius d1d-1.
Let $\bar{M}_{q}^{\star}(n;d)$ denote the maximum cardinality of a code in $\mathbb{Z}_{q}^{n}$ of minimum distance $d$ with respect to the distance function (4). As already noted at the beginning of the proof of Theorem 3.1, even though the distance function (4) is not a metric, the Gilbert–Varshamov bound still holds in the case analyzed in the present paper, implying that
$\bar{M}_{q}^{\star}(n;d)\;\geqslant\;\frac{q^{n}}{\frac{1}{q^{n}}\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};r)|}\;=\;\frac{q^{2n}}{\bar{T}_{q}(n;d-1)},$ (40)
where
$\bar{T}_{q}(n;r)\coloneqq\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};r)|$ (41)
is the “total ball size” of radius $r$. By Proposition 1, we have $M_{q}^{\star}(n;t)\geqslant\bar{M}_{q}^{\star}(n;2t)$, so (40) is automatically a lower bound on $M_{q}^{\star}(n;\lfloor\frac{d-1}{2}\rfloor)$.
Let $\bar{R}_{q}^{\star}(\delta)$ be the largest asymptotic rate of codes in $\mathbb{Z}_{q}^{n}$ of minimum distance $\delta n$, expressed in bits per symbol,
$\bar{R}_{q}^{\star}(\delta)\coloneqq\lim_{n\to\infty}\frac{1}{n}\log_{2}\bar{M}_{q}^{\star}(n;\delta n),$ (42)
and denote the exponential growth rate (to base 22) of $\bar{T}_{q}(n;\rho n)$ by
$\bar{\beta}_{q}(\rho)\coloneqq\lim_{n\to\infty}\frac{1}{n}\log_{2}\bar{T}_{q}(n;\rho n).$ (43)
Then (40) implies that
$\bar{R}^{\star}_{q}(\delta)\;\geqslant\;2\log_{2}(q)-\bar{\beta}_{q}(\delta).$ (44)
4.2 Estimating the average cardinality of $\bar{B}(\boldsymbol{x};r)$
Let $A(\boldsymbol{x};r)$ be the set of strings that can be obtained from $\boldsymbol{x}$ by exactly $r$ transpositions. Denote
$P_{q}(n;r)\coloneqq\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|A(\boldsymbol{x};r)|,$ (45)
and
$\alpha_{q}(\rho)\coloneqq\lim_{n\to\infty}\frac{1}{n}\log_{2}P_{q}(n;\rho n).$ (46)
Note that the maximum number of disjoint transpositions $r$ is $\frac{n}{2}$, so the domain of $\alpha_{q}(\cdot)$ is $0,\frac{1}{2}]$.
Lemma 5
For $q\geqslant 2$ and $\rho\in[0,\frac{1}{2}]$,
$\alpha_{q}(\rho)=(1-\rho)H\Big{(}\frac{\rho}{1-\rho}\Big{)}+(1-\rho)\log_{2}(q)+\rho\log_{2}(q-1),$ (47)
where $H(\cdot)$ is the binary entropy function. The function $\alpha_{q}(\rho)$ is concave and maximized for
$\rho^{*}=\frac{1}{2}\left(1-\sqrt{\frac{q}{5q-4}}\right).$ (48)
Proof
We shall derive the generating function of the bivariate sequence $(P_{q}(n;r))_{n,r}$ and use the methods of analytic combinatorics [24] to extract the exponent $\alpha_{q}(\rho)$ from it.
Consider the two cases:
  • $\bullet$
    $|A(\boldsymbol{x};r)|=|A(\boldsymbol{x}^{\prime}y;r)|+|A(\boldsymbol{x}^{\prime};r-1)|.$ (49a)
    The first term corresponds to the case when all the transpositions occur in the substring $\boldsymbol{x}^{\prime}y$, and the second term to the case when up to r1r-1 transpositions occur in the substring $\boldsymbol{x}^{\prime}$ and one transposition occurs in $yy^{\prime}$.
  • $\bullet$
    $|A(\boldsymbol{x};r)|=|A(\boldsymbol{x}^{\prime}y;r)|.$ (49b)
We now have
$F_{q}(z,u)$ $\coloneqq\sum_{n,r\geqslant 0}P_{q}(n;r)z^{n}u^{r}$
$=\sum_{n,r}\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|A(\boldsymbol{x};r)|z^{n}u^{r}$
$=1+\sum_{n\geqslant 1}\sum_{r\geqslant 0}\sum_{y\in\mathbb{Z}_{q}}\sum_{\boldsymbol{x}^{\prime}yy\in\mathbb{Z}_{q}^{n}}|A(\boldsymbol{x}^{\prime}y;r)|z^{n}u^{r}$
$\phantom{=\ 1}+\sum_{n\geqslant 1}\sum_{r\geqslant 0}\sum_{y\in\mathbb{Z}_{q}}\sum_{y^{\prime}\in\mathbb{Z}_{q}\setminus\{y\}}\sum_{\boldsymbol{x}^{\prime}yy^{\prime}\in\mathbb{Z}_{q}^{n}}(|A(\boldsymbol{x}^{\prime}y;r)|+|A(\boldsymbol{x}^{\prime};r-1)|)z^{n}u^{r}$
$=1+\sum_{n\geqslant 1}\sum_{r\geqslant 0}\sum_{\boldsymbol{x}^{\prime}y\in\mathbb{Z}_{q}^{n-1}}|A(\boldsymbol{x}^{\prime}y;r)|z^{n}u^{r}$
$\phantom{=\ 1}+(q-1)\sum_{n\geqslant 1}\sum_{r\geqslant 0}\sum_{\boldsymbol{x}^{\prime}y\in\mathbb{Z}_{q}^{n-1}}|A(\boldsymbol{x}^{\prime}y;r)|z^{n}u^{r}$
$\phantom{=\ 1}+q(q-1)\sum_{n\geqslant 1}\sum_{r\geqslant 0}\sum_{\boldsymbol{x}^{\prime}\in\mathbb{Z}_{q}^{n-2}}|A(\boldsymbol{x}^{\prime};r-1)|z^{n}u^{r}$
$=1+zF_{q}(z,u)+(q-1)zF_{q}(z,u)+q(q-1)z^{2}uF_{q}(z,u)$
$=1+(qz+q(q-1)z^{2}u)F_{q}(z,u),$ (50)
and hence
$F_{q}(z,u)=\frac{1}{1-qz-q(q-1)z^{2}u}.$ (51)
Denote by $G_{q}(z,u)=1-qz-q(q-1)z^{2}u$ the denominator in (51). We now use [23, Theorem 1.3], which implies that the exponential growth rate of $P_{q}(n;\rho n)$ equals
$\alpha_{q}(\rho)=-\log_{2}(z^{*})-\rho\log_{2}(u^{*}),$ (52)
where $z^{*},u^{*}$ are the unique positive solutions to the equations $G_{q}(z,u)=0$ and $\rho z\frac{\partial G_{q}(z,u)}{\partial z}=u\frac{\partial G_{q}(z,u)}{\partial u}$, i.e.,
$1-qz-q(q-1)z^{2}u$ $=0$ (53a)
$\rho(qz+2q(q-1)z^{2}u)$ $=q(q-1)z^{2}u.$ (53b)
Solving these equations, we obtain
$z^{*}=\frac{1-2\rho}{q(1-\rho)},\quad u^{*}=\frac{q\rho(1-\rho)}{(q-1)(1-2\rho)^{2}},$ (54)
and plugging the solutions back into (52) gives
$\alpha_{q}(\rho)=$ $-\rho\log_{2}(\rho)-(1-2\rho)\log_{2}(1-2\rho)+(1-\rho)\log_{2}(1-\rho)$
$+(1-\rho)\log_{2}(q)+\rho\log_{2}(q-1)$
$=$ $\ (1-\rho)H\Big{(}\frac{\rho}{1-\rho}\Big{)}+(1-\rho)\log_{2}(q)+\rho\log_{2}(q-1).$ (55)
The second part of the statement is shown by calculating the derivatives of the function $\alpha_{q}(\rho)$. $\blacksquare$
Recall that $B(\boldsymbol{x};r)$ represents the set of strings that can be obtained from $\boldsymbol{x}$ by at most $r$ disjoint transpositions. Denote
$T_{q}(n;r)\coloneqq\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|B(\boldsymbol{x};r)|$ (56)
and
$\beta_{q}(\rho)\coloneqq\lim_{n\to\infty}\frac{1}{n}\log_{2}T_{q}(n;\rho n).$ (57)
Lemma 6
For $q\geqslant 2$ and $\rho\geqslant 0$, we have
$\beta_{q}(\rho)=\max_{\rho^{\prime}\in[0,\rho]}\alpha_{q}(\rho^{\prime})=\begin{cases}\alpha_{q}(\rho),&\rho\leqslant\rho^{*}\\ \alpha_{q}(\rho^{*}),&\rho\geqslant\rho^{*}\end{cases}$ (58)
where $\alpha_{q}(\cdot)$ and $\rho^{*}$ are given in Lemma 5.
Proof
Since $B(\boldsymbol{x};r)=\bigcup_{r^{\prime}=0}^{r}A(\boldsymbol{x};r^{\prime})$, we have
$\max_{r^{\prime}\in\{0,\ldots,r\}}|A(\boldsymbol{x};r^{\prime})|\;\leqslant\;|B(\boldsymbol{x};r)|\;\leqslant\;\sum_{r^{\prime}=0}^{r}|A(\boldsymbol{x};r^{\prime})|$ (59)
and hence
$\max_{r^{\prime}\in\{0,\ldots,r\}}P_{q}(n;r^{\prime})\;\leqslant\;T_{q}(n;r)\;\leqslant\;\sum_{r^{\prime}=0}^{r}P_{q}(n;r^{\prime})\;\leqslant\;(r+1)\max_{r^{\prime}\in\{0,\ldots,r\}}P_{q}(n;r^{\prime}).$ (60)
It follows from (60) that $\beta_{q}(\rho)=\max_{\rho^{\prime}\in[0,\rho]}\alpha_{q}(\rho^{\prime})$, which is, by the fact that $\alpha_{q}(\rho)$ is concave and maximized for $\rho=\rho^{*}$, equal to the right-hand side of (58). $\blacksquare$
We now use the above result to obtain an upper bound on $\bar{\beta}_{q}(\rho)$. Note that, by Lemma 1, the domain of this function is $0,1]\cup\{\infty\}$.
Lemma 7
For any $q\geqslant 2$ and $\rho\in[0,1]$,
$\bar{\beta}_{q}(\rho)\;\leqslant\;\max_{\lambda\in[\frac{\rho}{3},\rho]}\left[2\lambda H\Big{(}\frac{\rho-\lambda}{2\lambda}\Big{)}+\beta_{q}(\lambda)\right],$ (61)
where $\beta_{q}(\cdot)$ is given in Lemma 6.
Proof
Lemma 4 implies that
$\bar{T}_{q}(n;r)$ $\;\leqslant\;\Big{(}\frac{2r}{3}+1\Big{)}\max_{v\in\{0,\ldots,2r/3\}}\binom{2(r-v)}{v}T_{q}(n;r-v)$
$\;=\;\Big{(}\frac{2r}{3}+1\Big{)}\max_{l\in\{r/3,\ldots,r\}}\binom{2l}{r-l}T_{q}(n;l).$ (62)
This, together with the fact that $\lim_{n\to\infty}\frac{1}{n}\log_{2}\!\binom{\gamma n}{\theta n}=\gamma H(\frac{\theta}{\gamma})$, gives (61). $\blacksquare$
4.3 Lower bound on $\bar{R}^{\star}_{q}(\delta)$
The following statement gives a lower bound on the achievable asymptotic rates of codes of relative distance $\delta$ (with respect to the distance function (4)).
Theorem 4.1
For any $q\geqslant 2$ and $\delta\in[0,1]$,
$\bar{R}^{\star}_{q}(\delta)\;\geqslant\;2\log_{2}(q)-\max_{\lambda\in[\frac{\delta}{3},\delta]}\left[2\lambda H\Big{(}\frac{\delta-\lambda}{2\lambda}\Big{)}+\beta_{q}(\lambda)\right],$ (63)
where the function $\beta_{q}(\cdot)$ is given in Lemma 6.
Proof
The result follows from the generalized Gilbert–Varshamov bound (44) and the upper bound on $\bar{\beta}_{q}(\cdot)$ given in Lemma 7. $\blacksquare$
Remark 3 (Lower bound for the binary case)
For $q=2$, a better lower bound on $\bar{R}^{\star}_{2}(\delta)$ is known [9, Section IV.C], as is the case in the regime of constant $t$ (see Remark 2). This is again due to the fact that the binary case seems to be rather special and admits a geometric interpretation in terms of packings in the Manhattan metric. $\blacklozenge$
5 Codes correcting all patterns of transpositions
It turns out that exponentially large codes of minimum distance $\infty$ exist in the model studied in this paper. In other words, the zero-error capacity of the transposition channel is positive, for any $q\geqslant 2$. For the special case $q=2$, it was shown in [21, Corollary 20] that $\bar{R}^{\star}_{2}(\infty)\geqslant\log_{2}(\lambda_{1})\approx 0.587$, where $\lambda_{1}$ is the largest root of the polynomial $p_{1}(x)=x^{7}-3x^{4}-2$. This was improved in [2, Theorem 3] to
$\bar{R}^{\star}_{2}(\infty)\;\geqslant\;\log_{2}(\lambda_{2})\;\approx\;0.643,$ (64)
where $\lambda_{2}\approx 1.561$ is the largest root of the polynomial $p_{2}(x)=x^{6}-2x^{3}-2x^{2}-2$. For $q$ divisible by 33, it was shown in [21, Theorem 21] that $\bar{R}^{\star}_{q}(\infty)\geqslant\log_{2}(\frac{q}{3})$. A construction that is valid for all $q$ and that improves upon [21, Theorem 21] was given in [2, Section IV], showing that
$\bar{R}^{\star}_{q}(\infty)\;\geqslant\;\frac{1}{2}\log_{2}\!\left(\left\lfloor\frac{q}{2}\right\rfloor\!\cdot\!\left\lceil\frac{q}{2}\right\rceil\right).$ (65)
In this section we provide a lower bound on $\bar{R}^{\star}_{q}(\infty)$ that is valid for arbitrary $q$, and that coincides with (64) for $q=2$ and improves upon (65) for $q\in\{3,4\}$. It is based on a code construction that represents a generalization of the construction from [2, Section III] for the binary case.
Note that any lower bound on $\bar{R}^{\star}_{q}(\infty)$ is automatically a lower bound on $\bar{R}^{\star}_{q}(\delta)$ for all $\delta\in[0,1]$, and is potentially better than (63) for $\delta>\delta_{0}$ (see Figure 1). In other words, for any $\delta\in[0,1]$, we can write
$\bar{R}^{\star}_{q}(\delta)\;\geqslant\;\max\!\left\{R^{\textsc{gv}}_{q}(\delta),\bar{R}^{\star}_{q}(\infty)\right\},$ (66)
where $R^{\textsc{gv}}_{q}(\delta)$ is the expression on the right-hand side of (63).
@@ 5 Codes correcting all patterns of transpositions
$\mathcal{P}\coloneqq\big{\{}(a,a,a),\,$ $(b,b,b),(a,b,b,b),(b,a,a,a),$
$(a,a,b,b,b,b),(b,b,a,a,a,a)\colon a\in\mathcal{A}_{0},b\in\mathcal{A}_{1}\big{\}}.$ (67)
@@
For every $q\geqslant 2$ and $n\geqslant 1$, we have $d_{\min}(\mathcal{D}_{q}(n))=\infty$, i.e., the code $\mathcal{D}_{q}(n)$ corrects all possible patterns of transpositions. Consequently,
$\bar{R}^{\star}_{q}(\infty)\;\geqslant\;\log_{2}(\lambda_{q}),$ (68)
where $\lambda_{q}$ is the unique positive root of the polynomial $p_{q}(x)=x^{6}-qx^{3}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil x^{2}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil$.
@@
For the first part of the statement, we need to show that an arbitrary codeword $\boldsymbol{x}\in\mathcal{D}_{q}(n)$ can be unambiguously recovered from the channel output $\boldsymbol{y}$ obtained after arbitrarily many transpositions have been applied to $\boldsymbol{x}$. Let $\boldsymbol{\tilde{x}}=(\tilde{x}_{1},\ldots,\tilde{x}_{n})$ denote the binary indicator string for $\boldsymbol{x}$ with respect to the partition $\{\mathcal{A}_{0},\mathcal{A}_{1}\}$, meaning that $\tilde{x}_{i}=j$ if and only if $x_{i}\in\mathcal{A}_{j}$, $j\in\{0,1\}$, and let $\widetilde{\mathcal{D}}(n)\coloneqq\{\boldsymbol{\tilde{x}}\colon\boldsymbol{x}\in\mathcal{D}_{q}(n)\}$. It was shown in [2, Theorem 3] that $\widetilde{\mathcal{D}}(n)$ is a binary code correcting all patterns of transpositions. We now have the following situation: the receiver can infer $\boldsymbol{\tilde{y}}$ from $\boldsymbol{y}$ (because it knows the sets $\mathcal{A}_{0},\mathcal{A}_{1}$), and then infer $\boldsymbol{\tilde{x}}$ from $\boldsymbol{\tilde{y}}$ (because $\boldsymbol{\tilde{x}}$ is a codeword of a transposition-correcting code), and what is left for it to do is infer $\boldsymbol{x}$ from $\boldsymbol{\tilde{x}}$. It is easy to see, by analyzing the possible concatenations of blocks from (67), that this can always be done. E.g., $(a,a,a,a^{\prime},a^{\prime},a^{\prime})$ may have resulted in $(a,a,a^{\prime},a,a^{\prime},a^{\prime})$ in the channel, $(a,a,a,a^{\prime},b,b,b)$ may have resulted in $(a,a,a^{\prime},a,b,b,b)$, etc. (here $a,a^{\prime}\in\mathcal{A}_{0}$, $b\in\mathcal{A}_{1}$), but such errors can easily be recognized and corrected because $(a,a,a^{\prime})$ is not a valid block. This means that the decoding process $\boldsymbol{y}\to\boldsymbol{\tilde{y}}\to\boldsymbol{\tilde{x}}\to\boldsymbol{x}$ is unambiguous, and hence the code $\mathcal{D}_{q}(n)$ corrects all patterns of transpositions.
It follows from the construction that the code size satisfies the recurrence relation
$|\mathcal{D}_{q}(n)|=$ $|\mathcal{A}_{0}|\cdot|\mathcal{D}_{q}(n-3)|+|\mathcal{A}_{1}|\cdot|\mathcal{D}_{q}(n-3)|$
$+2|\mathcal{A}_{0}||\mathcal{A}_{1}|\cdot|\mathcal{D}_{q}(n-4)|+2|\mathcal{A}_{0}||\mathcal{A}_{1}|\cdot|\mathcal{D}_{q}(n-6)|.$ (69)
This implies that $|\mathcal{D}_{q}(n)|=\lambda_{q}^{n+o(n)}$ as $n\to\infty$, where $\lambda_{q}$ is the unique positive root of the characteristic polynomial $p_{q}(x)=x^{6}-(|\mathcal{A}_{0}|+|\mathcal{A}_{1}|)x^{3}-2|\mathcal{A}_{0}||\mathcal{A}_{1}|x^{2}-2|\mathcal{A}_{0}||\mathcal{A}_{1}|$. Taking $|\mathcal{A}_{0}|=\lfloor\frac{q}{2}\rfloor$, $|\mathcal{A}_{1}|=\lceil\frac{q}{2}\rceil$, the second part of the statement follows. $\blacksquare$
@@
Using (66), (65), and Theorems 4.1 and 5.1, we have, for any $q\geqslant 2$ and any $\delta\in[0,1]$,
$\bar{R}^{\star}_{q}(\delta)\;\geqslant\;\max\!\left\{R^{\textsc{gv}}_{q}(\delta),\,\log_{2}(\lambda_{q}),\,\frac{1}{2}\log_{2}\!\left(\left\lfloor\frac{q}{2}\right\rfloor\!\cdot\!\left\lceil\frac{q}{2}\right\rceil\right)\right\},$ (70)
where $R^{\textsc{gv}}_{q}(\delta)$ is the expression on the right-hand side of (63), and $\lambda_{q}$ is the unique positive root of the polynomial $p_{q}(x)=x^{6}-qx^{3}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil x^{2}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil$. The expression $\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil$ simplifies to $\frac{q^{2}}{4}$ for even $q$, and $\frac{q^{2}-1}{4}$ for odd $q$.

### v2

@@ 3.1 Generalized Gilbert–Varshamov bound
For $q=2$, a lower bound on ${R}^{\star}_{2}(\tau)$ better than (30) is known [12, Section IV.C]. This is due to the fact that binary codes correcting transposition (a.k.a. bit-shift) errors can be interpreted as packings in the Manhattan metric [21], which in particular leads to a finer estimate of $\bar{T}_{2}(n;r)$. Unfortunately, the mentioned geometric interpretation fails when $q>2$ and thus the corresponding analysis and the constructions via dense packings cannot be generalized to larger alphabets in any obvious way. $\blacklozenge$
3.2 Codes correcting an arbitrary number of transpositions and the zero-error bound
It turns out that exponentially large codes of minimum distance $\infty$ exist in the model studied in this paper. In other words, the zero-error capacity of the transposition channel – the largest asymptotic rate of codes correcting an arbitrary number of transpositions – is positive, for any $q\geqslant 2$. Note that, by (7), Proposition 1(a), and (26), this rate is precisely ${R}^{\star}_{q}(1/2)$.
For the special case $q=2$, it was shown in [26, Corollary 20] that ${R}^{\star}_{2}(1/2)\geqslant\log_{2}(\xi_{1})\approx 0.587$, where $\xi_{1}$ is the largest root of the polynomial $p_{1}(x)=x^{7}-3x^{4}-2$. This was improved in [4, Theorem 3] to
${R}^{\star}_{2}(1/2)\;\geqslant\;\log_{2}(\xi_{2})\;\approx\;0.643,$ (39)
where $\xi_{2}\approx 1.561$ is the largest root of the polynomial $p_{2}(x)=x^{6}-2x^{3}-2x^{2}-2$. For $q$ divisible by 33, it was shown in [26, Theorem 21] that ${R}^{\star}_{q}(1/2)\geqslant\log_{2}(q/3)$. A construction that is valid for all $q$ and that improves upon [26, Theorem 21] was given in [4, Section IV], showing that
${R}^{\star}_{q}(1/2)\;\geqslant\;\log_{2}\!\sqrt{\lfloor q/2\rfloor\!\cdot\!\lceil q/2\rceil}.$ (40)
In this section we provide a lower bound on ${R}^{\star}_{q}(1/2)$ that is valid for arbitrary $q$, and that coincides with (39) for $q=2$ and improves upon (40) for $q\in\{3,4\}$. It is based on a code construction that represents a generalization of the construction from [4, Section III] for the binary case.
Note that any lower bound on ${R}^{\star}_{q}(1/2)$ is automatically a lower bound on ${R}^{\star}_{q}(\tau)$ for all $\tau\in[0,\frac{1}{2}]$, and is potentially better than (30) for $\tau>\tau_{0}$ (see Figure 1). In other words, for any $\tau\in[0,\frac{1}{2}]$, we can write
${R}^{\star}_{q}(\tau)\;\geqslant\;\max\!\left\{R^{\textsc{gv}}_{q}(\tau),\,{R}^{\star}_{q}(1/2)\right\},$ (41)
where $R^{\textsc{gv}}_{q}(\tau)$ is the expression on the right-hand side of (30).
@@ 3.2 Codes correcting an arbitrary number of transpositions and the zero-error bound
$\mathcal{P}\coloneqq\big\{(a,a,a),$ $(b,b,b),(a,b,b,b),(b,a,a,a),$ (42)
$(a,a,b,b,b,b),(b,b,a,a,a,a)\colon a\in\mathcal{A}_{0},b\in\mathcal{A}_{1}\big\}.$
@@
For every $q\geqslant 2$ and $n\geqslant 1$, we have $d_{\min}(\mathcal{D}_{q}(n))=\infty$, i.e., the code $\mathcal{D}_{q}(n)$ corrects an arbitrary number of transpositions. Consequently,
${R}^{\star}_{q}(1/2)\;\geqslant\;\log_{2}(\xi_{q}),$ (43)
where $\xi_{q}$ is the unique positive root of the polynomial $p_{q}(x)=x^{6}-qx^{3}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil x^{2}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil$.
@@
For the first part of the statement, we need to show that any codeword $\boldsymbol{x}\in\mathcal{D}_{q}(n)$ can be unambiguously recovered from the channel output $\boldsymbol{y}$ obtained after arbitrarily many transpositions have been applied to $\boldsymbol{x}$. Let $\boldsymbol{\tilde{x}}=(\tilde{x}_{1},\ldots,\tilde{x}_{n})$ denote the binary indicator string for $\boldsymbol{x}$ with respect to the partition $\{\mathcal{A}_{0},\mathcal{A}_{1}\}$, meaning that $\tilde{x}_{i}=j$ if and only if $x_{i}\in\mathcal{A}_{j}$, $j\in\{0,1\}$, and let $\widetilde{\mathcal{D}}(n)\coloneqq\{\boldsymbol{\tilde{x}}\colon\boldsymbol{x}\in\mathcal{D}_{q}(n)\}$. It was shown in [4, Theorem 3] that $\widetilde{\mathcal{D}}(n)$ is a binary code correcting an arbitrary number of transpositions. We then have the following situation: the receiver can infer $\boldsymbol{\tilde{y}}$ from $\boldsymbol{y}$ because it knows the sets $\mathcal{A}_{0},\mathcal{A}_{1}$, and it can infer $\boldsymbol{\tilde{x}}$ from $\boldsymbol{\tilde{y}}$ because $\boldsymbol{\tilde{x}}$ is a codeword of a transposition-correcting code. Hence, what is left for it to do is infer $\boldsymbol{x}$ from the pair $\boldsymbol{\tilde{x}}$, $\boldsymbol{y}$. It is easy to see, by analyzing the possible concatenations of blocks from (42), that this can always be done. First, by using $\boldsymbol{\tilde{x}}$ as the indicator string, all the transpositions in $\boldsymbol{y}$ that involve symbols from different sets, $a\in\mathcal{A}_{0}$, $b\in\mathcal{A}_{1}$, can be corrected. Second, the transpositions involving symbols from the same set, e.g., $a,a^{\prime}\in\mathcal{A}_{0}$, can also be easily recognized and corrected because they result in invalid blocks. For example, $(a,a,a,a^{\prime},a^{\prime},a^{\prime})$ may have resulted in $(a,a,a^{\prime},a,a^{\prime},a^{\prime})$ in the channel, $(a,a,a,a^{\prime},b,b,b)$ may have resulted in $(a,a,a^{\prime},a,b,b,b)$, etc. This means that the decoding process $\boldsymbol{y}\to(\boldsymbol{y},\boldsymbol{\tilde{y}})\to(\boldsymbol{y},\boldsymbol{\tilde{x}})\to\boldsymbol{x}$ is unambiguous, and hence the code $\mathcal{D}_{q}(n)$ corrects all patterns of transpositions.
It follows from the construction that the code cardinality satisfies the recurrence relation
$|\mathcal{D}_{q}(n)|=$ $|\mathcal{A}_{0}|\cdot|\mathcal{D}_{q}(n-3)|+|\mathcal{A}_{1}|\cdot|\mathcal{D}_{q}(n-3)|$ (44)
$+2|\mathcal{A}_{0}||\mathcal{A}_{1}|\cdot|\mathcal{D}_{q}(n-4)|+2|\mathcal{A}_{0}||\mathcal{A}_{1}|\cdot|\mathcal{D}_{q}(n-6)|.$
This implies that $|\mathcal{D}_{q}(n)|=\xi_{q}^{n+o(n)}$ as $n\to\infty$, where $\xi_{q}$ is the unique positive root of the characteristic polynomial $p_{q}(x)=x^{6}-(|\mathcal{A}_{0}|+|\mathcal{A}_{1}|)x^{3}-2|\mathcal{A}_{0}||\mathcal{A}_{1}|x^{2}-2|\mathcal{A}_{0}||\mathcal{A}_{1}|$. Taking $|\mathcal{A}_{0}|=\lfloor\frac{q}{2}\rfloor$, $|\mathcal{A}_{1}|=\lceil\frac{q}{2}\rceil$, the second part of the statement follows. $\blacksquare$
@@
Using (41), (40), and Theorems 3.1 and 3.2, we have, for any $q\geqslant 2$ and any $\tau\in[0,\frac{1}{2}]$,
${R}^{\star}_{q}(\tau)\;\geqslant\;\max\!\left\{R^{\textsc{gv}}_{q}(\tau),\,\log_{2}(\xi_{q}),\,\log_{2}\!\sqrt{\lfloor q/2\rfloor\!\cdot\!\lceil q/2\rceil}\right\},$ (45)
where $R^{\textsc{gv}}_{q}(\tau)$ is the expression on the right-hand side of (30), and $\xi_{q}$ is the unique positive root of the polynomial $p_{q}(x)=x^{6}-qx^{3}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil x^{2}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil$. The expression $\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil$ simplifies to $\frac{q^{2}}{4}$ for even $q$, and $\frac{q^{2}-1}{4}$ for odd $q$.